# Exploring Adversarial Obstacle Attacks in Search-based Path Planning for Autonomous Mobile Robots

### v1

@@ I INTRODUCTION
While these technological advances have been significant, the focus has primarily been on functionality and innovations, leaving safety and security considerations, particularly in robotics, considerably behind. The environments where the robots are being deployed, such as logistics, public spaces and streets, and disaster areas, are rarely threat-free and can be subject to attacks from adversaries. While some studies have been conducted, the vulnerability of mobile robots to adversarial attacks remains largely underexplored. The need for cyber-security and cyber-safety in robotics differs from securing traditional computing systems due to the added dimension of physical interaction [1]. This added aspect means that gaining control over a robot could result in harm, not only to the robot itself but also to its immediate environment, posing a potential and serious risk of injury to nearby individuals.
@@
Given the threats faced by autonomous mobile robots in their operational environments, it is critical to develop robust path planning algorithms that can withstand adversarial manipulations. This study examines the robustness of search-based A* path planning against adversarial manipulations in the operational environment. It explores the implications for safety and security in autonomous robotics using a methodologically simple yet effective ‘obstacle attack’. While previous studies [2, 3, 4] have focused on completely immobilising or severely restricting the agent’s movement, we focus on causing delays in the agent’s path-finding efforts, thereby reducing its likelihood of detection by defensive mechanisms. This is achieved by determining obstacles’ optimal placement and subsequent integration into the operational environment to delay the robot’s timely arrival at its destination. We conducted a set of experiments, both simulation within a static warehouse setting and on a physical robot in small environments, and collected data to demonstrate the effectiveness of the attacks. We studied the time required for the robotic agent to reach its goal and the strategic placement of obstacles.
@@ II RELATED WORK
Path planning has been heavily studied in the past [5, 6, 7, 8]. Here, we briefly review the work related to adversarial attacks in robots.
@@
While there have been some studies on the security threats that affect robotic hardware, firmware/OS, and applications [9, 10, 11, 12], less focus has been given to studying the vulnerabilities of AI deployed in robotics [13]. Robots interacting with a physical environment bring a new category of threats beyond software and hardware risks referred to as behaviour threats [14]. For example, an autonomous delivery robot can make nearby pedestrians feel threatened, prompting them to take action to render the robot immobile [15].
@@
Adversarial attacks can significantly impact autonomous robots, disrupting control systems and decision-making capabilities. Vemprala et al. [2] investigate the susceptibility of iterative optimisation-based trajectory planners to adversarial attacks, showing that such attacks can disrupt the optimisation process by manipulating the cost function and lead to failures or increased computation times. The authors test their method by applying adversarial strategies to two state-of-the-art trajectory planners, demonstrating that these planners are consistently vulnerable to such attacks. Cavorsi et al. [3] present an algorithm for resilient path planning in robot teams, where robots can coordinate despite adversaries. The authors use Control Barrier Functions (CBFs) to balance resilience and safe navigation, developing a method to identify and handle trade-offs between these objectives. Wu et al. [4] focus on the vulnerability of robot motion planning algorithms to adversarial attacks, specifically in the physical environment. Their study highlights two types of attacks: planner failure and blindspot attacks. Planner failure attacks involve making subtle changes to the physical environment, causing the motion planner to fail to find a valid path. In contrast, blindspot attacks exploit occlusions and sensor field-of-view limitations, leading the planner to generate a collision-free trajectory that intersects with unperceived obstacles. The reported experimental results indicate that these attacks can drastically increase failure rates to $95\%$ and collision rates to $90\%$ with only minor environmental modifications.
@@
Clark et al. [16] examine the manipulation of an autonomous robotic vehicle’s ultrasonic collision avoidance sensors to redirect its path. By employing the Q-learning algorithm for real-time route control, they illustrate the potential for indirect attacks on robotic systems, emphasising the need for robust countermeasures. Shi et al. [17] explore the vulnerabilities of learning-based locomotion controllers in quadrupedal robots using adversarial attacks. Their study focuses on the robustness of controllers that employ deep Reinforcement Learning (RL) against real-world uncertainties like sensor noise and external perturbations. Despite empirical robustness, this research shows that these controllers can fail significantly under carefully crafted, low-magnitude adversarial sequences.
@@
Lechner et al. [18] studied the robustness-accuracy trade-off, exploring the balance between maintaining accuracy and ensuring robustness against adversarial attacks. Focusing on two neural controllers - one for vision-based commands and another for LiDAR-based motor control; their findings suggest that adversarial training can introduce new errors, potentially undermining the robot’s learned behaviour. Lechner et al. [19] revisit the robustness-accuracy trade-off, incorporating advancements like overparametrisation and vision transformers. Their study shows that integrating adversarial weight perturbation reduces the gap between robustness and accuracy.
@@ III-A Threat Model
Knowledge & Capability: We assume that the adversary could carry out the attack virtually by manipulating the agent’s cost map to create a false obstacle. This manipulation could be achieved through a man-in-the-middle attack [20] or malicious software. To calculate the optimal obstacle placement, the adversary would need access to the map of the environment in which the agent operates. We assume that the adversary possesses this knowledge, as well as the agent’s starting (current) and goal positions. These assumptions are reasonable, given that robots repeatedly operate in the same environment.
@@ III-B Obstacle Attack
1:  $path\leftarrow\textsc{A*}(start,goal,map)$
2:  $path_{longest}\leftarrow path$
3:  for $step$ in $path$ do
4:     $path_{new}\leftarrow$ run A* with obstacle at $step$
@@
6:        $obstacle_{best}\leftarrow$ obstacle at $step$
@@ III-C2 Real-world Deployment
A schematic representation of the real-world deployment of the obstacle attack is depicted in Fig. 3. We validate the performance of the attack in real-world conditions on a Unitree Go1. To leverage the existing robot navigation frameworks, the attack is integrated with the 2D ROS navigation stack to act as a part of the global path planner. Given the cost map, the attack generates an obstacle in the global path, making sure the obstacle does not fully block all the possible paths. The local planner and obstacle avoidance algorithm used is the Timed-Elastic-Band (TEB) [22, 23] with localisation provided by the Phasespace tracking system in real-time at $960$ Hz. Examples of successful deployment of the pipeline can be seen in Fig. 1.
@@ IV RESULTS
Scenario Map Mean absolute delay Mean percentage delay Success rate

### v2

@@ I INTRODUCTION
While these technological advances have been significant, the focus has primarily been on functionality and innovations, leaving safety and security considerations, particularly in robotics, considerably behind. The environments where the robots are being deployed, such as logistics, public spaces and streets, and disaster areas, are rarely threat-free and can be subject to attacks from adversaries. While some studies have been conducted, the vulnerability of mobile robots to adversarial attacks remains largely underexplored. The need for cyber-security and cyber-safety in robotics differs from securing traditional computing systems due to the added dimension of physical interaction [Portugal2017]. This added aspect means that gaining control over a robot could result in harm, not only to the robot itself but also to its immediate environment, posing a potential and serious risk of injury to nearby individuals.
@@
Given the threats faced by autonomous mobile robots in their operational environments, it is critical to develop robust path planning algorithms that can withstand adversarial manipulations. This study examines the robustness of search-based A* path planning against adversarial manipulations in the operational environment. It explores the implications for safety and security in autonomous robotics using a methodologically simple yet effective ‘obstacle attack’. While previous studies [vemprala2021adversarial, cavorsi2023multi, Wu2024] have focused on completely immobilising or severely restricting the agent’s movement, we focus on causing delays in the agent’s path-finding efforts, thereby reducing its likelihood of detection by defensive mechanisms. This is achieved by determining obstacles’ optimal placement and subsequent integration into the operational environment to delay the robot’s timely arrival at its destination. We conducted a set of experiments, both simulation within a static warehouse setting and on a physical robot in small environments, and collected data to demonstrate the effectiveness of the attacks. We studied the time required for the robotic agent to reach its goal and the strategic placement of obstacles.
@@ II RELATED WORK
Path planning has been heavily studied in the past [ellis2022navigation, liu2023vit, liu2024dipper, stamatopoulou2024dippest]. Here, we briefly review the work related to adversarial attacks in robots.
@@
While there have been some studies on the security threats that affect robotic hardware, firmware/OS, and applications [Clark2017, Creado2017, AhmadYousef2018, Yaacoub2021], less focus has been given to studying the vulnerabilities of AI deployed in robotics [Calo2018]. Robots interacting with a physical environment bring a new category of threats beyond software and hardware risks referred to as behaviour threats [Colledanchise2021]. For example, an autonomous delivery robot can make nearby pedestrians feel threatened, prompting them to take action to render the robot immobile [Oravec2022].
@@
Adversarial attacks can significantly impact autonomous robots, disrupting control systems and decision-making capabilities. Vemprala et al. [vemprala2021adversarial] investigate the susceptibility of iterative optimisation-based trajectory planners to adversarial attacks, showing that such attacks can disrupt the optimisation process by manipulating the cost function and lead to failures or increased computation times. The authors test their method by applying adversarial strategies to two state-of-the-art trajectory planners, demonstrating that these planners are consistently vulnerable to such attacks. Cavorsi et al. [cavorsi2023multi] present an algorithm for resilient path planning in robot teams, where robots can coordinate despite adversaries. The authors use Control Barrier Functions (CBFs) to balance resilience and safe navigation, developing a method to identify and handle trade-offs between these objectives. Wu et al. [Wu2024] focus on the vulnerability of robot motion planning algorithms to adversarial attacks, specifically in the physical environment. Their study highlights two types of attacks: planner failure and blindspot attacks. Planner failure attacks involve making subtle changes to the physical environment, causing the motion planner to fail to find a valid path. In contrast, blindspot attacks exploit occlusions and sensor field-of-view limitations, leading the planner to generate a collision-free trajectory that intersects with unperceived obstacles. The reported experimental results indicate that these attacks can drastically increase failure rates to $95\%$ and collision rates to $90\%$ with only minor environmental modifications.
@@
Clark et al. [Clark2018] examine the manipulation of an autonomous robotic vehicle’s ultrasonic collision avoidance sensors to redirect its path. By employing the Q-learning algorithm for real-time route control, they illustrate the potential for indirect attacks on robotic systems, emphasising the need for robust countermeasures. Shi et al. [Shi2024] explore the vulnerabilities of learning-based locomotion controllers in quadrupedal robots using adversarial attacks. Their study focuses on the robustness of controllers that employ deep Reinforcement Learning (RL) against real-world uncertainties like sensor noise and external perturbations. Despite empirical robustness, this research shows that these controllers can fail significantly under carefully crafted, low-magnitude adversarial sequences.
@@
Lechner et al. [Lechner2021] studied the robustness-accuracy trade-off, exploring the balance between maintaining accuracy and ensuring robustness against adversarial attacks. Focusing on two neural controllers - one for vision-based commands and another for LiDAR-based motor control; their findings suggest that adversarial training can introduce new errors, potentially undermining the robot’s learned behaviour. Lechner et al. [Lechner2023] revisit the robustness-accuracy trade-off, incorporating advancements like overparametrisation and vision transformers. Their study shows that integrating adversarial weight perturbation reduces the gap between robustness and accuracy.
@@ III-A Threat Model
Knowledge & Capability: We assume that the adversary could carry out the attack virtually by manipulating the agent’s cost map to create a false obstacle. This manipulation could be achieved through a man-in-the-middle attack [conti2016survey] or malicious software. To calculate the optimal obstacle placement, the adversary would need access to the map of the environment in which the agent operates. We assume that the adversary possesses this knowledge, as well as the agent’s starting (current) and goal positions. These assumptions are reasonable, given that robots repeatedly operate in the same environment.
@@ III-B Obstacle Attack
[…]
3:for stepstep in pathpath do
4:  $path_{new}\leftarrow$ run A* with obstacle at stepstep
@@
6:   $obstacle_{best}\leftarrow$ obstacle at stepstep
@@ III-C2 Real-world Deployment
A schematic representation of the real-world deployment of the obstacle attack is depicted in Fig. 3. We validate the performance of the attack in real-world conditions on a Unitree Go1. To leverage the existing robot navigation frameworks, the attack is integrated with the 2D ROS navigation stack to act as a part of the global path planner. Given the cost map, the attack generates an obstacle in the global path, making sure the obstacle does not fully block all the possible paths. The local planner and obstacle avoidance algorithm used is the Timed-Elastic-Band (TEB) [rosmann2012trajectory, rosmann2013efficient] with localisation provided by the Phasespace tracking system in real-time at $960$ Hz. Examples of successful deployment of the pipeline can be seen in Fig. 1.
@@ IV RESULTS
Scenario Map
Mean
absolute
delay
Mean
percentage
delay
Success
rate